# Combined Fit of Spectrum and Composition for FR0 Radio Galaxy Emitted Ultra-High-Energy Cosmic Rays with Resulting Secondary Photons and Neutrinos

### v1

@@ 4.2 Emitted Nuclei
Combined fits to the energy spectrum and estimated mean log mass number () of data are done using sets of emitted proton, helium, nitrogen, silicon, and iron nuclei. These nuclei are commonly chosen in combined fit analyses as each elemental species can be considered as a family of neighboring nuclei and this selection represents a comprehensive sampling of cosmic ray nuclei observed in composition spectra. Studies suggest that heavier nuclei can originate from specific types of astrophysical environments, such as the dense, magnetized surroundings of AGN and elements heavier than iron should be rare enough to have a negligible impact on combined fits.
The emitted nuclei propagation is modeled according to their interaction cross-section with the photon fields, energy loss due to radiation emission, and deflection by cosmic magnetic fields as described in Section 3, and are modified by the time they are observed. The resulting combined fit in bins of observed energy is shown in Figure 5 for an example configuration and illustrates how different nuclei are modified from emission to detection. The observed slightly varies between the fitted configurations depending on the energy spectrum fit parameters spectral index , rigidity-dependent exponential cutoff Rcut, and trajectory cutoff Dcut. Emitted proton is, of course, unmodified and has a . The emitted nuclei resulting observed for all evolving fraction fits (see Section 5.2) is shown in A.
@@ 5 Combined Fit of Composition and Energy Spectrum
5.1 Constant Fraction Fits
For the constant energy-independent observed fraction fit, the simulated FR0 UHECR emission that best matches Auger data [27, 28] is found by minimizing the summed per degree of freedom for the composition and energy spectrum, considering eight input parameters. This fitting process utilizes four parameters representing the fractions of emitted protons, helium, nitrogen, silicon, and iron. These fractions, spanning the entire energy spectrum, are constrained to sum to 100%. The remaining parameters include the emitted spectral index , rigidity-dependent exponential cutoff R [37, 38, 39], maximum particle trajectory cutoff D, and energy spectrum normalization .
The total cost function minimized is the sum of the per degree-of-freedom for composition and energy spectrum comparisons between simulation and data:
@@ 5.1 Constant Fraction Fits
5.2 Evolving Fraction Fit
For the evolving energy-dependent observed fraction fit, the simulated FR0 UHECR emission that best matches Auger data [27, 28] is found by minimizing the summed per bin for the composition and energy spectrum. This approach involves forty-eight input parameters, providing a detailed, bin-specific analysis. Forty-four parameters represent the fractions of emitted protons, helium, nitrogen, silicon, and iron, calculated for each of the eleven energy bins used in the Auger composition data, with each bin’s total fraction constrained to 100%. The remaining parameters are the same energy spectrum parameters as Section 5.1. Though an evolving nuclei fraction dependent on energy is not a good model discriminator due to over-fitting, it is more capable of reflecting physical processes than a constant observed fraction model. This detailed modeling provides a deeper picture of the emission requirements for each EAS model and magnetic field for the best possible fit.
The cost function minimized is the sum of the per bin for composition and energy spectrum comparisons between simulation and data and is minimized first using the constant fraction fit as an initial search value. The fitting process then iteratively refines the model by splitting element fractions into increasingly detailed energy ranges (two energy ranges, then four fractions, and finally eleven fractions, one for each energy bin), applying deterministic nested fitting at each step. As before the nested fitting holds the spectrum parameters constant while varying the elemental content found in the previous step then vice versa. Additional improvement of the result is done by a nested fit one energy bin at a time (keeping all others constant) and then a final fit of all 48 parameters at once.
As before, this method facilitates an extensive global search of possible values of the parameters, though given the results of the constant fraction fit can be more constrained to without a loss in bootstrap uncertainty calculation. Uncertainties on the fit parameters are calculated the same as Section 5.1.
@@ 6 Results
6.1 Constant Fraction Fit
The constant observed fraction combined fit results for all five simulated intergalactic media and two EAS models (EPOS-LHC and Sibyll2.3c) compared to Auger data [27, 28] are shown in Figure 6. Results for the QGSJetII-04 model are detailed in subsequent tables. For the observed mean log mass number (Figure 6a), the 1nG random fields show a better fit for the lower energy, lighter composition, bins. In contrast, only the Dolag and Rand.B models sufficiently account for the data in the highest energy bin of the energy spectrum, as shown in Figure 6b — though the Dolag model overshoots the previous two bins. The excellent fit of the longer correlation length 1 nG (Rand.B) field at the highest energies is somewhat unexpected, if FR0s do not contribute significantly at these extreme energies. The CLUES model, being on the order of J(E)*E3 1033 eV2km-2yr-1str-1, essentially does not contribute to the highest energy bin.
Table 1 and Table 3 detail the best-fit parameters for each model configuration, along with the corresponding 1 equivalent confidence intervals that provide a measure of the uncertainty around these estimates. Using an 8-dimensional multivariate (all fit parameters) kernel density estimation (KDE), Table 2 and Table 4 present the most probable values and 1 standard deviations derived from the bootstrap sample distributions.
Table 1 and Table 2 show each EAS model and magnetic field’s total , emission spectral index (), exponential rigidity-dependent cutoff (log10()), particle trajectory cutoff (), and relative energy spectrum normalization (). Table 3 and Table 4 show the fitted FR0 emission elemental fractions.
@@ 6.1 Constant Fraction Fit
Models with minimal magnetic fields, such as the ’no-field’ or Dolag scenarios, are disfavored for all three EAS models, indicating a significant influence of magnetic fields on fit accuracy. Notably, the QGSJetII-04 EAS model typically shows a much poorer fit quality compared to the others, except when combined with the random fields, suggesting a sensitivity to specific intergalactic conditions.
6.1.1 Energy Spectrum Parameters
Analysis of the constant fraction energy spectrum parameters of Table 1 and Table 2 suggest that both the magnetic field strength B and RMS BRMS influence the expected cosmic-ray source emission energy spectrum. However, the results are perhaps surprisingly stable given the wide range of magnetic field models and EAS model compositions. As B or BRMS increases, the emission spectral index does decrease – with generally harder emission spectra for QGSJetII-04. The fit values of the exponential rigidity-dependent cutoff, log10(), generally increases with both B and BRMS. For smaller magnetic fields, the rigidity cutoff decreases with a heavier elemental composition, while it increases with larger field strengths. The trajectory cutoff fit values, , remain fairly consistent across the fields tested, with the exception of the Dolag field, where a significantly higher BRMS of 11 nG introduces a noticeable deviation.
Also, for any given magnetic field, EAS models that result in a heavier composition (Sibyll2.3c is heaviest as shown in Figure 6a and QGSJetII-04 is the lightest) leads to a softer (increasing) spectral index E-γ. An exception occurs with the bootstrap KDE most probable for Dolag-SIBYLL shown in Table 2, as this configuration fit leads to multi-modal distributions (a 1-d KDE fit results in =2.30).
Compared to the previous FR0 combined fit results, with sources limited to z 0.05 (and a slightly older CRPropa version), the emission spectra for the QGSJetII-04 and EPOS-LHC EAS models are harder (a smaller ). The exponential cutoff has also increased for all scenarios, with the exception of the Rand.B magnetic field [24].
@@ 6.1.1 Energy Spectrum Parameters
Excluding the disfavored magnetic fields (no-field and Dolag) and the disfavored CLUES-QGS4 configuration, the bounds found for the best fit of these parameters from Table 1 are:
1.
1.97 2.54
2.
19.45 log10(/V) 19.86
3.
842 Mpc Dcut 854 Mpc
4.
1.341 1.354
@@
For the bootstrap most probable parameters from Table 2, excluding the disfavored configurations, the ranges found are:
1.
@@
2.
19.41 log10(/V) 19.87
3.
842 Mpc Dcut 864 Mpc
4.
@@
Compared to the recent Auger result of [37] these parameter values are similar to the reported second minima of their models and the low z source distance model (z0.02).
@@ 6.1.2 UHECR Composition Parameters
The FR0 emission fractions for constant (observed) fraction fits of emitted proton, helium, nitrogen, silicon, and iron nuclei are given in Table 3, Table 4, and Figure 7. Due to significant propagation effects, where a large percentage of emitted helium transforms into observed protons (as detailed in Figure 5), the best fit emission fractions for protons and helium exhibit some degeneracy. This results in instability in determining whether protons or helium dominate for a given magnetic field. The combined light component fraction () and the corresponding KDE most probable values of Table 4 show more stable results.
Table 3, Table 4, and Figure 7 display the emission fractions for protons, helium, nitrogen, silicon, and iron as part of the constant observed fraction fits of FR0 emitted nuclei. These results indicate that magnetic field strength can significantly impact the relative nuclei abundances of the source-emitted cosmic rays. As the magnetic field strength increases, the emission fractions of protons and helium remains relatively stable. In contrast, the fractions of nitrogen nuclei tend to increase, whereas heavier nuclei such as silicon and iron show a decrease in their relative fractions.
Compared to the previous FR0 combined fit results, with sources limited to z 0.05, the helium fractions have significantly increased for the QGSJetII-04 and EPOS-LHC EAS models (except for the no-field scenario) [24]. The amount of proton has commensurately decreased for these configurations (except for the no-field scenario). Also, comparatively, there is now more nitrogen for the Rand.A-EPOS and Rand.B-EPOS configurations and less nitrogen for the Rand.A-QGS4 and Rand.B-QGS4 configurations. The fraction of silicon has decreased for all configurations except for a small increase for Rand.B-QGS4. Finally, the iron fraction has increased for all magnetic fields except for the no-field scenario [24].
Excluding the disfavored configurations, the ranges found for the FR0 emission best fit elemental fractions (in %) from Table 3 are:
1.
58.4 98.9
2.
0.0 74.5
3.
10.1 69.3
4.
0.0 38.8
5.
0.0 1.3
6.
1.1 8.2
The analysis shows a broad range of elemental fractions, with the most notable being a substantial variation in the combined proton and helium fraction, ranging from 58.4% to 98.9%, and the fraction of nitrogen ranging from zero to 38.8%. These ranges reflect the effect of interactions between magnetic fields and nuclei during cosmic ray propagation.
For the bootstrap most probable parameters from Table 4, excluding the disfavored configurations, the ranges found for the FR0 emission elemental fractions (%) are:
1.
59.3 98.4
2.
5.3 81.2
3.
9.5 63.9
4.
0 35.8
5.
0 0.7
6.
0.7 7.5
@@
What may be expected from the constant fraction combined fits using the older Dominguez11 IR model is an increase in emitted helium and a decrease in proton with a hard emission spectrum – though this result could simply be due to fitting uncertainties [37, 38]. A more thorough comparison done in [46] with soft injection iron nuclei shows a smaller variation on the energy spectrum and a maximum 10% difference on at 1019.6 eV.
6.1.3 Multi-messenger Photons and Neutrinos
@@
The constant fraction fit cosmogenic integral photon and all-flavor neutrino spectra are shown in Figures 8a and 8b, respectively, for the best-fit configuration (compared to the no-field scenario) – the CLUES structured magnetic field and Sibyll2.3c-based mass composition. Although the Sibyll2.3c hadronic interaction model results in a heavier reconstructed mass for data compared to the other models, it predicts a 92% emission of light nuclei (protons and helium) from FR0 sources. At lower energies, the integral photon flux aligns closely with a theoretical prediction for a pure proton cosmogenic source, while at higher energies, it transitions into a range predicted for a mixed composition.
@@
Notably, the neutrino flux resembles predictions for a pure iron source, a result that may be influenced by the simulation constraint to relatively proximate FR0 sources (z0.2). However, although the previous FR0 combined fit results, limited to sources with z0.05, have fewer low-energy neutrinos and more high-energy neutrinos, the differences are not significant [24]. Considering the broad spectrum of theoretical predictions and current experimental upper limits, the simulated FR0 fluxes falls within a reasonable range. Overall, in both photon and neutrino spectra, the presence of a magnetic field generally leads to increased fluxes, with this effect being most pronounced at the highest energies, as shown in Figure 9.
6.2 Evolving Fraction Fits

### v2

@@ 4.2 Emitted Nuclei
Combined fits to the energy spectrum and estimated mean log mass number () of data are done using sets of emitted proton, helium, nitrogen, silicon, and iron nuclei. These nuclei are commonly chosen in combined-fit analyses as each elemental species can be considered as a family of neighboring nuclei and this selection represents a comprehensive sampling of cosmic-ray nuclei observed in composition spectra. Studies suggest that heavier nuclei can originate from specific types of astrophysical environments, such as the dense, magnetized surroundings of AGN and elements heavier than iron should be rare enough to have a negligible impact on combined fits.
The emitted nuclei propagation is modeled according to their interaction cross-section with the photon fields, energy loss due to radiation emission, and deflection by cosmic magnetic fields as described in Section 3, and are modified by the time they are observed. The resulting combined fit in bins of observed energy is shown in Figure 5 for an example configuration and illustrates how different nuclei are modified from emission to detection. The observed slightly varies between the fitted configurations depending on the energy-spectrum fit parameters spectral index , rigidity-dependent exponential cutoff R, and trajectory cutoff D. Emitted proton is, of course, unmodified and has a . The emitted nuclei resulting observed for all evolving-fraction fits (see Section 5.2) is shown in Appendix A.
@@ 5 Combined Fit of Composition and Energy Spectrum
5.1 Constant-Fraction Fits
For the constant energy-independent observed fraction fit, the simulated FR0 UHECR emission that best matches Auger data [27, 28] is found by minimizing the summed per degree of freedom for the composition and energy spectrum, considering eight input parameters. This fitting process utilizes four parameters representing the fractions of emitted protons, helium, nitrogen, silicon, and iron. These fractions, spanning the entire energy spectrum, are constrained to sum to 100%. The remaining parameters include the emitted spectral index , rigidity-dependent exponential cutoff R [37, 38, 39], maximum particle-trajectory cutoff D, and the energy-spectrum-only normalization .
The total cost function minimized is the sum of the per degree-of-freedom for composition and energy-spectrum comparisons between simulation and data:
The degrees of freedom for the energy spectrum are taken as and for the composition .
@@ 5.1 Constant-Fraction Fits
5.2 Evolving-Fraction Fit
For the evolving energy-dependent observed fraction fit, the simulated FR0 UHECR emission that best matches Auger data [27, 28] is found by minimizing the summed per bin for the composition and energy spectrum. This approach involves forty-eight input parameters, providing a detailed, bin-specific analysis. Forty-four parameters represent the fractions of emitted protons, helium, nitrogen, silicon, and iron, calculated for each of the eleven energy bins used in the Auger composition data, with each bin’s total fraction constrained to 100%. The remaining parameters are those described in Section 5.1, namely the emitted spectral index , rigidity-dependent exponential cutoff R [37, 38, 39], maximum particle trajectory cutoff D, and the energy-spectrum normalization .
A constant fraction of emitted nuclei over the energy range has been assumed in previous UHECR combined-fit analyses such as discussed in Section 5.1. However, this assumption is unphysical, as the pre-acceleration composition of nuclei does not remain constant with energy. Additionally, acceleration processes, spallation products, and energy losses are all composition-dependent, further complicating the assumption of constancy across the energy spectrum. Though an evolving nuclei fraction dependent on energy is not a good model discriminator due to over-fitting, it is more capable of reflecting physical processes than a constant observed fraction model. This detailed modeling provides a deeper picture of how the best possible fit emission requirements vary with energy for each EAS model and magnetic field.
With essentially no degrees of freedom for composition across 44 nuclei fraction parameters, the cost function minimized is simply the sum of the chi-squared values per bin for the composition and energy-spectrum comparisons between simulation and data. The minimization begins using the constant-fraction fit as an initial search value. The fitting process then iteratively refines the model by splitting the elemental fractions into progressively more detailed energy ranges—first two ranges, then four, and finally eleven, corresponding to one fraction per energy bin. At each step, deterministic nested fitting is applied, holding the spectrum parameters constant while varying the elemental composition obtained from the previous step, and then vice versa. Additional refinement is achieved by performing a nested fit, varying one energy bin at a time while keeping the others constant, followed by a final fit of all 48 parameters simultaneously.
As before, this method facilitates an extensive global search of possible values of the parameters, though given the results of the constant-fraction fit can be more constrained to without a loss in bootstrap uncertainty calculation. Uncertainties on the fit parameters are calculated the same as Section 5.1.
@@ 6 Results
6.1 Constant-Fraction Fit
The constant observed fraction combined-fit results for all five simulated intergalactic media and two EAS models (EPOS-LHC and Sibyll2.3c) compared to Auger data [27, 28] are shown in Figure 6. Results for the QGSJetII-04 model are detailed in subsequent tables. For the observed mean log mass number (Figure 6(a)), the 1 nG random fields show a better fit for the lower energy, lighter composition, bins. In contrast, only the Dolag and Rand.B models sufficiently account for the data in the highest energy bin of the energy spectrum, as shown in Figure 6(b)—though the Dolag model overshoots the previous two bins. The excellent fit of the longer correlation length 1 nG (Rand.B) field at the highest energies is somewhat unexpected, if FR0s do not contribute significantly at these extreme energies. The CLUES model, being on the order of J(E)*E3 1033 eV2km-2yr-1str-1, essentially does not contribute to the highest energy bin.
Table 2 and Table 4 detail the best-fit parameters for each model configuration, along with the corresponding 1 equivalent confidence intervals that provide a measure of the uncertainty around these estimates. Using an 8-dimensional multivariate (all fit parameters) kernel density estimation (KDE), Table 2 and Table 4 present the most-probable values and 1 standard deviations derived from the bootstrap sample distributions.
Table 2 and Table 2 show each EAS model and magnetic field’s total , emission spectral index (), exponential rigidity-dependent cutoff (log10(R/V)), particle trajectory cutoff (D), and relative energy-spectrum normalization (). Table 4 and Table 4 show the fitted FR0 emission elemental fractions.
@@ 6.1 Constant-Fraction Fit
Configurations with minimal magnetic fields, such as the ’no-field’ or Dolag scenarios, are disfavored for all three EAS models as they have 23, indicating a significant influence of magnetic fields on fit accuracy. Notably, the QGSJetII-04 EAS model typically shows a much poorer fit quality compared to the others, except when combined with the random fields, suggesting a sensitivity to specific intergalactic conditions.
6.1.1 Energy-Spectrum Parameters
@@
Analysis of the constant-fraction energy-spectrum parameters of Table 2 and Table 2 suggest that both the magnetic field strength B and RMS BRMS influence the expected cosmic-ray source emission energy spectrum. However, the results are perhaps surprisingly stable given the wide range of magnetic field models and EAS model compositions. As B or BRMS increases, the emission spectral index does decrease—with generally harder emission spectra for QGSJetII-04. The fit values of the exponential rigidity-dependent cutoff, R, generally increases with both B and BRMS. For smaller magnetic fields, R decreases with a heavier elemental composition, while it increases with larger field strengths. The trajectory cutoff fit values, D, remain fairly consistent across the fields tested, with the exception of the Dolag field, where a significantly higher BRMS of 11 nG introduces a noticeable deviation.
Also, for any given magnetic field, EAS models that result in a heavier composition (Sibyll2.3c is heaviest as shown in Figure 6(a) and QGSJetII-04 is the lightest) leads to a softer (increasing) spectral index E-γ. An exception occurs with the bootstrap KDE most-probable for Dolag-SIBYLL shown in Table 2, as this configuration fit leads to multi-modal distributions (a 1-d KDE fit results in = 2.30).
Compared to the previous FR0 combined-fit results, with sources limited to 0.05 (and a slightly older CRPropa version), the emission spectra for the QGSJetII-04 and EPOS-LHC EAS models are harder (a smaller ). The exponential cutoff has also increased for all scenarios, with the exception of the Rand.B magnetic field [24].
Excluding the disfavored magnetic fields (no-field and Dolag) and the disfavored CLUES-QGS4 configuration (those with 23), the bounds found for the best fit of these parameters from Table 2 are:
@@
19.45 log10(R/V) 19.86
842 D/Mpc 854
1.341 1.354
For the bootstrap most-probable parameters from Table 2, excluding the disfavored configurations, the ranges found are:
2.15 2.54
19.41 log10(R/V) 19.87
842 D/Mpc 864
@@
Compared to the recent Auger result of [37] these parameter values are similar to the reported second minima of their models and the low- source distance model ( 0.02).
@@ 6.1.2 UHECR Composition Parameters
The FR0 emission fractions for constant (observed) fraction fits of emitted proton, helium, nitrogen, silicon, and iron nuclei are shown in Figure 7, Table 4, and Table 4. Due to significant propagation effects, where a large percentage of emitted helium transforms into observed protons (as detailed in Figure 5), the best-fit emission fractions for protons and helium exhibit some degeneracy. This results in instability in determining whether protons or helium dominate for a given magnetic field. The combined light component fraction () and the corresponding KDE most-probable values of Table 4 show more stable results.
These results indicate that magnetic field strength can significantly impact the relative nuclei abundances of the source-emitted cosmic rays. As the magnetic field strength increases, the emission fractions of protons and helium remains relatively stable. In contrast, the fractions of nitrogen nuclei tend to increase, whereas heavier nuclei such as silicon and iron show a decrease in their relative fractions.
Compared to the previous FR0 combined-fit results, with sources limited to 0.05, the helium fractions have significantly increased for the QGSJetII-04 and EPOS-LHC EAS models (except for the no-field scenario) [24]. The amount of proton has commensurately decreased for these configurations (except for the no-field scenario). Also, comparatively, there is now more nitrogen for the Rand.A-EPOS and Rand.B-EPOS configurations and less nitrogen for the Rand.A-QGS4 and Rand.B-QGS4 configurations. The fraction of silicon has decreased for all configurations except for a small increase for Rand.B-QGS4. Finally, the iron fraction has increased for all magnetic fields except for the no-field scenario [24].
Excluding the disfavored configurations, the ranges found for the FR0 emission best-fit elemental fractions (in %) from Table 4 are:
58.4 98.9
0.0 74.5
10.1 69.3
0.0 38.8
0.0 1.3
1.1 8.2
The analysis shows a broad range of elemental fractions, with the most notable being a substantial variation in the combined proton and helium fraction, ranging from 58.4% to 98.9%, and the fraction of nitrogen ranging from zero to 38.8%. These ranges reflect the effect of interactions between magnetic fields and nuclei during cosmic-ray propagation.
For the bootstrap most-probable parameters from Table 4, excluding the disfavored configurations, the ranges found for the FR0 emission elemental fractions (%) are:
59.3 98.4
5.3 81.2
9.5 63.9
0 35.8
0 0.7
0.7 7.5
@@
What may be expected from the constant-fraction combined fits using the older Dominguez11 IR model is an increase in emitted helium and a decrease in proton with a hard emission spectrum—though this result could simply be due to fitting uncertainties [37, 38]. A more thorough comparison done in [47] with soft injection iron nuclei shows a smaller variation on the energy spectrum and a maximum 10% difference on at an energy of E = 1019.6 eV.
6.1.3 Multimessenger Photons and Neutrinos
@@
The constant-fraction fit cosmogenic integral photon and all-flavor neutrino spectra are shown in Figures 8(a) and 8(b), respectively, for the best-fit configuration (compared to the no-field scenario)—the CLUES structured magnetic field and Sibyll2.3c-based mass-composition. Although the Sibyll2.3c hadronic interaction model results in a heavier reconstructed mass for data compared to the other models, it predicts a 92% emission of light nuclei (protons and helium) from FR0 sources propagated through the CLUES magnetic field. At lower energies, the integral photon flux aligns closely with a theoretical prediction for a pure proton cosmogenic source, while at higher energies, it transitions into a range predicted for a mixed composition.
@@
Notably, the neutrino flux resembles predictions for a pure iron source, a result possibly influenced by the simulation’s constraint to relatively nearby FR0 sources ( 0.2). However, although the previous FR0 combined-fit results, limited to sources with 0.05, show fewer low-energy neutrinos and more high-energy neutrinos, the differences are not significant [24]. Considering the broad range of theoretical predictions and current experimental upper limits, the simulated FR0 fluxes falls within a reasonable range. Overall, in both photon and neutrino spectra, the presence of a magnetic field generally leads to increased fluxes, with this effect being most pronounced at the highest energies, as shown in Figure 9.
The total number of cosmogenic photons largely depends on the fraction of emitted protons. Looking at the total number of photons with energies E1016 eV, shown by the first bin of Figure 9(a), the configuration with the highest photon count is the No-Field-EPOS which has the highest proton emission at 93.5%. The lowest photons count is from Rand.A-SIBYLL, which has the lowest proton emission at 0% (the next two lowest photon counts come from the Dolag configurations with 0% proton emission). Without a magnetic field, the number of detected photons is approximately 55 times lower for iron than for protons, as the energy is divided among the 56 nucleons, reducing the energy available for photopion production. Additionally, given the same emission, an intergalactic magnetic field increases the number of detected photons. A proton emission with = 1, without exponential cutoff or trajectory cutoff (just the raw simulation before fitting), results in approximately three times more photons in the Rand.B 1ng-3Mpc magnetic field than in the no-field case.
As with photons, the number of detected neutrinos increases with the magnetic field. Comparing no-field to the Rand.B 1ng-3Mpc case, protons produce 1.03 times more neutrinos, while iron produces 1.6 times more with the magnetic field. However, inversely from photons, the number of detected neutrinos increases with mass number: the no-field iron raw simulation results in approximately 1.3 times more neutrinos than the proton case.
In summary, all things being equal for a given source emission, turning on an intergalactic magnetic field increases the number of cosmogenic photons and neutrinos. Since we are fitting to data, each configuration’s emission is different, leading to differences in composition and energy spectra. The energy spectrum is also important, as can most easily be seen in the no-field scenario, which has the softest emission spectrum and strongest exponential cutoff, resulting in the lowest photon and neutrino counts at the highest energies.
6.2 Evolving-Fraction Fits